### v1

# LogitSpec: Accelerating Retrieval-based Speculative Decoding via Next Next Token Speculation

Abstract
Speculative decoding (SD), where a small draft model is employed to propose draft tokens in advance and then the target model validates them in parallel, has emerged as a promising technique for LLM inference acceleration. Many endeavors to improve SD are to eliminate the need for a draft model and generate draft tokens in a retrieval-based manner in order to further alleviate the drafting overhead and significantly reduce the difficulty in deployment and applications. However, retrieval-based SD relies on a matching paradigm to retrieval the most relevant reference as the draft tokens, where these methods often fail to find matched and accurate draft tokens. To address this challenge, we propose LogitSpec to effectively expand the retrieval range and find the most relevant reference as drafts. Our LogitSpec is motivated by the observation that the logit of the last token can not only predict the next token, but also speculate the next next token. Specifically, LogitSpec generates draft tokens in two steps: (1) utilizing the last logit to speculate the next next token; (2) retrieving relevant reference for both the next token and the next next token. LogitSpec is training-free and plug-and-play, which can be easily integrated into existing LLM inference frameworks. Extensive experiments on a wide range of text generation benchmarks demonstrate that LogitSpec can achieve up to 2.61$\times$ speedup and 3.28 mean accepted tokens per decoding step. Our code is available at https://github.com/smart-lty/LogitSpec.
1 Introduction
Large Language Models (LLMs), such as GPT-4.5 (OpenAI, 2024), DeepSeek R1 (DeepSeek-AI, 2025), Qwen 2.5 (Team, 2025), and LLaMA-3 (Team, 2024), have demonstrated remarkable capabilities across a wide range of natural language processing tasks, including question answering (Calijorne Soares and Parreiras, 2020), code generation (Jiang et al., 2024), and dialogue systems (Yi et al., 2024). While they achieve success by scaling up the model parameters, the increase in scale comes with significant inference challenges. The most straightforward challenge is auto-regressive decoding, where each LLM decoding step can only generate one token. This token-by-token decoding process incurs exacerbated latency with both the length of generated tokens and the model scale.
To address this challenge, speculative decoding (SD) (Leviathan et al., 2023; Chen et al., 2023) has emerged as a promising approach for lossless LLM inference acceleration. The key idea of SD is to employ a small draft model to first generate $\gamma$ draft tokens, and then the target model validates these tokens in parallel within a single forward, ensuring the output distribution to be unchanged. However, deploying an extra draft model introduces considerable overhead and difficulties, including (a) complex implementation and deployment, especially in integration with other techniques,(Liu et al., 2025) (b) additional memory overhead, especially in long-context scenarios. Moreover, when there exists no available draft model for SD, it takes substantial training cost to construct a compact draft model (Li et al., 2025a). Therefore, many existing works dedicate to developing draft-model-free SD methods (Saxena, 2023; Fu et al., 2024; He et al., 2024), where the draft tokens are generated in a retrieval-based manner. At each decoding step, they retrieve the relevant reference according to the last few tokens and extract draft tokens from the reference. In this way, they can explicitly eliminate the need for a draft model and reduce the drafting overhead.
Figure 1: Illustration of vanilla retrieval-based SD method and our LogitSpec. (a) Retrieval-based SD retrievals the wrong token “triangle” according to the next token “the”. (b) LogitSpec first speculates the next next token “area”, and then retrievals the right relevant reference “area of” accord to “the area”. This simple example illustrates how LogitSpec utilizes the last logit to speculate the next next token and improves the retrieval accuracy.
Despite their promising efficacy and implementation simplicity, these methods rely on a matching paradigm to effectively retrieve the most relevant reference as draft tokens and often struggle to find matched and accurate tokens. As shown in Figure 1 (a), with a simple prompt “What is the area of the triangle?”, vanilla retrieval-based SD methods fail to effectively retrieve the right token “area” according to the next token “the”, as the most relevant reference is “the triangle”. Moreover, it is often the case that no matched reference can be found, e.g., there are no matched tokens in more than 30% of decoding steps in PLD (Saxena, 2023).
In this paper, we propose a simple yet effective retrieval-based SD framework, namely LogitSpec, which leverages the logit of the last token (last logit) to predict the next next token and improves the accuracy of the retrieval reference. LogitSpec is motivated by the observation that the last logit can speculate the next next token with a relatively high accuracy. Based on this observation, we use the speculated next next token as a guidance to retrieval reference. Specifically, LogitSpec generates draft tokens in two steps: (1) utilizing the last logit to speculate the next next token; (2) retrieving relevant reference for both the next token and the next next token. As shown in Figure 1 (b), with the speculated token “area”, LogitSpec successfully retrievals the correct draft tokens “area of”. This two-step process enables better retrieval accuracy—it can help filter the most relevant reference when the reference is redundant, while extending the searching space when there exists no relevant reference. To summarize, our contributions are:
  • (a)
    We empirically observe that the last logit can speculate the next next token with a relatively high accuracy. Unlike other retrieval techniques which are sensitive to the specific task, this property is robust and effective across various tasks.
  • (b)
    Building on this observation, we propose LogitSpec, a plug-and-play retrieval-based SD framework which can improve the retrieval accuracy and achieve better speedup.
  • (c)
    We conduct extensive experiments on various text generation benchmarks to demonstrate the effectiveness of LogitSpec. Notably, LogitSpec achieves up to 2.61$\times$ speedup and 3.28 mean accepted tokens per decoding step without the need for an extra draft model.
2 Related Work
In this section, we briefly review the two main directions of speculative decoding, draft-model-based speculative decoding and draft-model-free speculative decoding. More discussions are provided in Appendix A.
Draft-model-based speculative decoding.
This line of research focuses on employing an extra draft model for speculative decoding. Some existing works make their efforts on developing a more compact draft model with less drafting overhead. Medusa (Cai et al., 2024) pioneers this direction by adding extra MLP heads at the top of the target model, reusing the hidden states from the target model and predicting the next few tokens in parallel. However, the generation from these decoding heads is independent, which harms the accuracy of the draft tokens. Therefore, some works (Ankner et al., 2024; Xiao et al., 2024) propose to add sequential dependencies for better performance. Glide (Du et al., 2024) takes the similar idea of Medusa and propose to reuse KV cache from the target model. Recently, Eagle (Li et al., 2025b, 2024, a) has dominated this research line by training an auto-regressive head and scaling up the training data. However, while these methods achieve superior speedup, all of them rely on an extra draft model, which necessitate extra parameters or extensive training. The existence of the draft model significantly limits their application: (a) deploying a draft model requires complex and careful implementation. (b) Both the draft model and its KV cache require additional GPU memory, which may incur load unbalance. In the long context settings, even the KV cache for the draft model may exceed the capacity of 1 GPU and leads to significant resource competitions; (c) the training cost is substantial, and the draft model requires retraining when the target model is updated.
Draft-model-free speculative decoding.
This line of research focuses on eliminating the need for a draft model. Some existing works (Zhang et al., 2024a; Elhoushi et al., 2024; Xia et al., 2025) observe the layer sparsity of LLMs that not all the layers are necessary to predict the next token. They propose methods to generate draft tokens with a subset of the layers of target model. However, the number of skipped layers is limited, which affects the overall speedup. Recently, retrieval-based SD (He et al., 2024; Saxena, 2023; Fu et al., 2024) has dominated the draft-model-free speculative decoding by retrieving the relevant reference as the draft tokens. The relevant reference can be derived from the question and generated tokens, or external database. However, there exists two common issues during the retrieval process: (a) retrieval from the context often fails to find matched draft tokens; (b) retrieval from external database often struggles to find accurate draft tokens. These two issues motivate us to develop a more appropriate retrieval-based speculative decoding framework.
3 Background
Speculative Decoding.
Let $\bm{x}$ denote an input sequence (prefix), a speculative decoding step consists of a drafting phase and a verification phase. During the drafting phase, the draft model $\mathcal{M}_{q}$ is employed to give $\gamma$ draft tokens $x_{1},x_{2},...,x_{\gamma}$ by running $\gamma$ times model forward and sample. Here, we denote the output logit $\mathcal{M}_{q}(\bm{x}+[x_{1},...,x_{i-1}])$ as $q_{i-1}$, then each draft token is given by $x_{i}\sim q_{i-1},i=1,...,\gamma$. During the verification phase, the prefix $\bm{x}$ together with $\gamma$ draft tokens are sent to $\mathcal{M}_{p}$ for verification. The target model $\mathcal{M}_{p}$ inputs $\bm{x}+[x_{1},...,x_{\gamma}]$ and outputs the logits $p_{0},p_{1},...,p_{\gamma}$. Then SD sequentially verifies $x_{i}$ via speculative sampling (Leviathan et al., 2023; Chen et al., 2023), where the acceptance rate is given by:
$\alpha_{i}=\left\{\begin{aligned} &1\quad\qquad\quad p_{i-1}[x_{i}]\geq q_{i-1}[x_{i}],\\ &\frac{p_{i-1}[x_{i}]}{q_{i-1}[x_{i}]}\quad p_{i-1}[x_{i}]<q_{i-1}[x_{i}],\\ \end{aligned}\right.$ (1)
If SD rejects $x_{i}$, it will resample a token from $norm(\max(0,p_{i-1}-q_{i-1}))$, otherwise, SD accepts all the draft tokens and samples an additional token from $p_{\gamma}$. In this way, each SD step generates at least 1 token and at most $\gamma+1$, leading to theoretical lossless quality and efficiency acceleration.
Retrieval-based Speculative Decoding.
The retrieval-based SD methods generate draft tokens in a retrieval-based manner. As retrieval-based SD methods do not require an additional draft model, We omit the notation for $\mathcal{M}_{q}$ and abbreviate $\mathcal{M}_{p}$ as $\mathcal{M}$. A retrieval model $\mathcal{R}$ is employed to store $n$-grams (referred as reference) from corpus
=$\mathcal{R}=\{(x_{i}^{1},x_{i}^{2},...,x_{i}^{n})\}_{i=1}^{N}.$ (2)
Here, $N$ denotes the total number of $n$-grams in $\mathcal{R}$. Existing retrieval-based SD methods mainly differ in constructing $\mathcal{R}$. At each SD decoding step, given a query $m$-gram tokens $x_{q}^{1},x_{q}^{2},...,x_{q}^{m}$, the retrieval model first traversals the reference, finds matched $n$-grams and returns the subsequent tokens of the matched tokens:
$\textsc{Match}(\mathcal{R},(x_{q}^{1},x_{q}^{2},...,x_{q}^{m}))=\{(x_{i}^{m+1},...,x_{i}^{n})\}_{i\in\mathcal{S}},$ (3)
where $\mathcal{S}=\{i|x_{i}^{t}=x_{q}^{t},\forall t=1,...,m\}$. Intuitively, a larger $m$ leads to more precise matches but lower match probabilities.
4 Method
Figure 2: Motivated observations. (a) The last logit can speculate the next next token with a relatively high accuracy. In over 50% decoding steps, the next next token can be found in the top-60 entries within the last logit. This prediction ability exists consistently across different model sizes and architectures. (b) Compared with other retrieval-based methods, the prediction of last logit demonstrates superior robustness to the downstream task type. Both of these observations motivate us to utilize the last logit to guide the retrieval process.
In this section, we introduce our LogitSpec, a novel retrieval-based SD framework that utilizes last logit as a guidance for retrieval and effectively improves the retrieval performance. We first conduct a motivated experiment in Section 4.1, and then introduce the whole framework of LogitSpec in Section 4.2 and 4.3. An overview of LogitSpec is shown in Figure 3.
4.1 Motivated Observation
Drawing insights from multi-token prediction (MTP) (Gloeckle et al., 2024) that fine-tuned LLMs can predict multiple tokens in a single forward, we are interested in the ability of LLMs to predict the next next token without fine-tuning. We conduct a simple experiment to investigate the rank of the next next token in last logit, which is used to predict the next token. Specifically, we use a small fraction of Spec-Bench (Xia et al., 2024) that contains 13 sub-tasks (detailed in Appendix B) and 6 different LLMs as the backbone models. The experimental results, as shown in Figure 2, demonstrate that the last logit has a strong potential to predict the next next token. In over 50% decoding steps, the next next token can be found within the top-60 entries of the last logit. For Llama 3.1 8B (Team, 2024) with a large 128K vocabulary, this percentage even increases to 64% within top-60 entries of the last logit. The results of 6 different LLMs on 13 downstream tasks demonstrate that this observation holds consistently across different models (Vicuna, Llama 3.1 and Qwen) and scales (ranging from 0.5B to 33B).
We further conduct experiments on Spec-Bench with Vicuna 7B to investigate the robustness of the predictive capability of the last logit across different tasks. We propose a minimal speculative decoding algorithm with last logit, last logit decoding for this experiment, where we directly use the top-60 entries of the last logit to serve as the draft tokens. Each draft token is a guess to the next next token. If the target model accept any one of the draft tokens, the others are dropped. Consequently, at each decoding step, the accepted length is either 2 or 1. As shown in 2, this simple SD method yields mean accepted tokens per decoding step (MAT) of over 1.5. Compared with other retrieval-based SD methods such as PLD (Saxena, 2023) and REST (He et al., 2024), which are highly affected by the task type, the MAT of last logit decoding exhibits superior robustness to the task. These results demonstrate that the ability of the last logit to predict the next next token widely exists in modern LLMs themselves, and strongly motivate us to improve retrieval-based SD methods with last logit. We provide a detailed experimental setup for the motivation experiments in Appendix B.
Figure 3: An overview of LogitSpec. At each decoding step, LogitSpec first utilizes the top-$k$ entries of the last logit as the speculation to the next next token. Then, LogitSpec retrievals relevant references for both the next token and the next next token. Finally, LogitSpec organizes the draft tokens into a draft tree and prepares a tree attention for parallel verification.
4.2 LogitSpec Drafting
[…]
=$\mathcal{L}=\{\widetilde{x}_{i+2}\ |\ \widetilde{x}_{i+2}\in\textsc{Top}_{k}(p_{i})\land\widetilde{x}_{i+2}\neq x_{i+1}\}.$ (4)
Then, we use the retrieval model $\mathcal{R}$ to retrieval reference for both the next token and next next tokens:
$\mathcal{D}=\textsc{Match}(\mathcal{R},\overbrace{(x_{i-m+2},...,x_{i},x_{i+1})}^{m\text{-gram for next token}})\cup\{\textsc{Match}(\mathcal{R},\overbrace{(x_{i-m+3},...,x_{i},x_{i+1},\widetilde{x}_{i+2})}^{m\text{-gram for next nexttoken}})\}_{\widetilde{x}_{i+2}\in\mathcal{L}}$ (5)
Here, $m$-gram denotes the last $m$ tokens as the query to retrieval the matched reference. In our experiments, we retrieval draft tokens with $m=3$ first. If no matched tokens are found, we decrease $m$ to 2 and so on. For simplicity, we only use the user-input prompt and decoded tokens to construct the retrieval model $\mathcal{R}$, that is to say, our LogitSpec is query-independent, and will not be affected by other queries. We give the detailed implementation of retrieval process in Appendix D.1.
Taking the example in Figure 3, we first select the top-$k$ entries of the last logit as the speculation to the next next token. Then, we retrieval reference for both the next token “the” and the next next token “area”, and return “of the”. After that, we organize these draft tokens into a draft tree, where the next token serves as the root. Finally, token sequence “area of the” is accepted by $\mathcal{M}$.
After the retrieving process, we employ a simple pruning strategy to control the number of the draft tokens. Specifically, we do not prune the retrieval tokens for the next token. For each speculated next next token, we prune the retrieval tokens with a simple heuristic strategy: if the rank of the next next token is below than 8, we preserve 4 tokens; if the rank of the next next token is below than 32, we preserve 3 tokens; otherwise we only preserve the speculate next next token itself. The retrieval process is terminated until the total number of draft tokens exceeds a specific draft tree capacity ${K}$.
4.3 LogitSpec Verification
After retrieving multiple draft token sequences $\mathcal{D}$, we organize these token sequences into a draft tree and prepare a tree attention for parallel verification. Intuitively, this process involves flattening $\mathcal{D}$ into a draft token sequence $\mathbf{x}_{\text{draft}}$.
Then, we prepare an attention mask to make each draft token sequence invisible to other sequences. Let $\text{Tril}(l)$ denotes a causal mask with length $l$, and $m_{j}$ denotes the length of the $j$-th token sequence in $\mathcal{D}$, the attention mask is given by:
$A_{\text{draft}}=\operatorname{diag}(\text{Tril}(m_{1}),\text{Tril}(m_{2}),...,\text{Tril}(m_{j})).$ (6)
An illustration of this attention mask is shown in Figure 3(c) (note that the next token is visible to all sequences, corresponding to the first column). Both the preparation of the attention mask and the verification of the draft tree are consistent with previous works (Miao et al., 2024; Cai et al., 2024). We provide a pseudo code of the attention mask preparations in Appendix C.
5 Experiments
5.1 Experimental Setup
Tasks and Datasets.
We conduct experiments on various text generation benchmarks to evaluate the effectiveness of LogitSpec. Specifically, we first evaluate the performance of LogitSpec on Spec-Bench (Xia et al., 2024). Spec-Bench is a widely used comprehensive benchmark that covers diverse application scenarios, including multi-turn conversation (MT), translation (Trans), summarization (Sum), question answering (QA), mathematical reasoning (Math), retrieval-augmented generation (RAG). Furthermore, we evaluate the performance of LogitSpec on HumanEval (Chen et al., 2021), GSM8K (Cobbe et al., 2021) and CNN/DM (Nallapati et al., 2016), which are widely used text generation benchmarks in existing works.
Baselines.
In our main experiments, we compare our LogitSpec to existing state-of-the-art plug-and-play SD methods: (a) Lookahead Decoding (Fu et al., 2024): utilizing the dropped draft tokens to serve as the retrieval model; (b) REST (He et al., 2024): employing an external knowledge base as the retrieval model; (c) PLD (Saxena, 2023): extracting n-grams from the user-input prompts and decoded tokens as the retrieval model; and (d) vanilla SpS (Leviathan et al., 2023; Chen et al., 2023): employing a well trained small-size model Vicuna-68M as the draft model.
Implementation Details.
For all experiments, we follow the standard setting of Spec-Bench with Pytorch (Paszke et al., 2019) version of 2.6.0, transformers (Wolf et al., 2020) version of 4.37.1, and CUDA (NVIDIA et al., 2020) 12.4. The experiments where the model size is between 7B and 33B are conducted on single NVIDIA SXM A100 80G GPU, while the experiments for models with 70B size are conducted on 2 NVIDIA SXM A100 80G GPUs. The inference precision is set as float16. Following prior work, we conduct experiments with temperature 0 with batch size of 1. The maximal generation lengths are set to 1024. More implementation details can be found in Appendix D. In our experiments, we use two widely used metrics for evaluation: mean accepted tokens per decoding step (denoted as MAT) and overall speedup ratio (denoted as Speedup).
5.2 Main Results
We conduct extensive experiments on various text generation benchmarks to demonstrate the effectiveness of our LogitSpec. As presented in Tables 1 and 2, LogitSpec significantly outperforms retrieval-based SD baselines by a large margin. We further present the experiments of Llama 2 chat series models in Appendix E.
Table 1: Experimental results of LogitSpec on CNN/DM (Nallapati et al., 2016), GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) with Vicuna. We report the mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
Models Method CNN/DM GSM8K HumanEval Average
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Vicuna 7B Lookahead 1.54 1.28 1.92 1.64 1.79 1.57 1.75 1.50
REST 1.64 1.19 1.55 1.13 1.96 1.51 1.72 1.28
PLD 2.61 2.26 1.80 1.62 1.84 1.65 2.08 1.84
SpS 2.34 1.58 2.08 1.49 2.56 1.80 2.33 1.62
LogitSpec 3.28 2.61 2.69 2.20 2.62 2.24 2.86 2.35
Vicuna 13B Lookahead 1.46 1.12 1.88 1.61 1.75 1.57 1.70 1.43
REST 1.65 1.22 1.57 1.18 1.94 1.55 1.72 1.32
PLD 2.34 1.85 1.76 1.61 1.98 1.77 2.03 1.74
SpS 2.18 1.48 2.00 1.55 2.66 1.95 2.28 1.66
LogitSpec 2.90 2.17 2.59 2.13 2.88 2.47 2.79 2.26
Vicuna 33B Lookahead 1.50 1.18 1.89 1.44 1.66 1.33 1.68 1.32
REST 1.65 1.41 1.57 1.32 1.99 1.64 1.74 1.46
PLD 2.07 1.72 1.66 1.41 1.60 1.42 1.78 1.52
SpS 2.01 1.57 1.97 1.54 2.27 1.80 2.08 1.64
LogitSpec 2.66 2.01 2.46 1.93 2.41 1.95 2.51 1.96
Table 2: Experimental results of LogitSpec on Spec-Bench (Xia et al., 2024) with Vicuna. We report the speedup ratio on each sub task, mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
Models Method MT Trans Sum QA Math RAG MAT Speedup
Vicuna 7B Lookahead 1.40 1.14 1.19 1.24 1.55 1.09 1.66 1.27
REST 1.63 1.31 1.36 1.66 1.21 1.73 1.82 1.48
PLD 1.64 1.04 2.43 1.14 1.61 1.71 1.73 1.59
SpS 1.66 1.13 1.62 1.49 1.47 1.55 2.28 1.49
LogitSpec 1.92 1.39 2.68 1.70 2.22 1.87 2.44 2.01
Vicuna 13B Lookahead 1.30 1.06 1.20 1.12 1.48 1.12 1.63 1.22
REST 1.52 1.17 1.37 1.53 1.19 1.55 1.82 1.38
PLD 1.47 1.02 2.19 1.03 1.57 1.71 1.68 1.48
SpS 1.60 1.13 1.68 1.39 1.53 1.67 2.18 1.49
LogitSpec 1.89 1.43 2.33 1.43 2.23 1.93 2.32 1.93
Vicuna 33B Lookahead 1.32 1.08 1.20 1.06 1.54 1.15 1.61 1.24
REST 1.63 1.27 1.45 1.61 1.30 1.61 1.80 1.48
PLD 1.44 1.06 2.00 1.07 1.55 1.45 1.55 1.42
SpS 1.75 1.28 1.76 1.53 1.69 1.68 2.01 1.61
LogitSpec 1.77 1.38 2.15 1.37 2.00 1.69 2.13 1.76
[…]
5.3 Ablation Study
To further investigate the components of LogitSpec and provide more insights, we conduct extensive ablation studies on the aforementioned 4 benchmarks in Table 3. Specifically, we mainly focus on two components of LogitSpec: last logit and retrieval. We denote LogitSpec without the last logit decoding and only using retrieving as w/o last logit, and LogitSpec without retrieval and only using last logit decoding as w/o retrieval. As shown in Table 3, the absence of any component results in a performance degradation of the entire framework.
Our findings are as follows First, the absence of retrieval exhibits more importance to the final acceleration, which is consistent with the discussion in Section 4.2 that the theoretical upper bound of last logit decoding severely hinders its real-world application. Second, the absence of retrieval and the absence of last logit shows different affect in different sub-tasks. For example, the absence of retrieval decreases MAT of 0.87 on Spec-Bench with Vicuna 7B, while it decreases MAT of 1.71 on CNN/DM. In contrast, the absence of last logit leads to a more consistent MAT degradation across different tasks, which further highlights the robustness of last logit decoding. Finally, these results demonstrate that combining last logit with retrieval improve the retrieval performance and overall speedup.
[…]
Models Method Spec-Bench CNN/DM GSM8K HumanEval
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Vicuna 7B w/o last logit 1.72↓.72 1.27↓.74 2.66↓.62 2.24↓.37 1.81↓.88 1.63↓.57 1.89↓.73 1.69↓.55
w/o retrieval 1.57↓.87 1.24↓.77 1.57↓1.71 1.23↓1.38 1.71↓.98 1.39↓.81 1.69↓.93 1.39↓.85
LogitSpec 2.44 2.01 3.28 2.61 2.69 2.20 2.62 2.24
Vicuna 13B w/o last logit 1.66↓.66 1.23↓.70 2.34↓.56 1.81↓.36 1.78↓.81 1.61↓.52 2.02↓.86 1.75↓.72
w/o retrieval 1.59↓.73 1.19↓.74 1.58↓1.32 1.22↓.95 1.71↓.88 1.45↓.68 1.63↓1.25 1.37↓1.10
LogitSpec 2.32 1.93 2.90 2.17 2.59 2.13 2.88 2.47
Vicuna 33B w/o last logit 1.55↓.58 1.22↓.54 2.05↓.61 1.72↓.29 1.64↓.82 1.39↓.54 1.59↓.82 1.40↓.55
w/o retrieval 1.61↓.52 1.25↓.51 1.62↓1.04 1.26↓.75 1.69↓.77 1.33↓.60 1.60↓.81 1.34↓.61
LogitSpec 2.13 1.76 2.66 2.01 2.46 1.93 2.41 1.95
5.4 Case Study
Figure 4: Running time breakdown of the whole decoding process on Spec-Bench with Vicuna 7B.
In-depth Running Time Analysis.
To further investigate the effectiveness of LogitSpec, we conduct experiments to analyze the running time allocation within the whole decoding process. Specifically, there are five non-negligible components in LogitSpec, including (a)retrieving draft tokens: the process of retrieving reference for the next token and the next next tokens; (b) preparation: preparing attention mask for the draft tokens; (c) model forward: conducting one-pass model forward; (d) verification: validating the draft tokens with speculative sampling; (e) update: necessary update of KV cache and retrieval model. We present the running time breakdown in Figure 4.
As shown in Figure 4, model forward occupies the majority of wall-clock time. Compared with vanilla auto-regressive decoding, the overhead introduced by LogitSpec, i.e., retrieval overhead, only takes 1.17% of the whole decoding process. This overhead can be further alleviated by some parallel techniques, as the retrieval process of the next next token is independent. The process of verification and update takes 1.03% and 1.05% through the whole process, respectively, which brings negligible overhead as well.
Retrieval Performance.
As mentioned in Section 4.2, our LogitSpec can effectively expand the retrieval range and improve the retrieval performance. To validate this statement, we conduct a simple experiment to test the retrieval successful rate, i.e., whether the model successfully retrievals reference as draft tokens. As shown in Table 4, PLD cannot retrieval any matched tokens in more than 30% decoding steps, while LogitSpec can retrieval matched reference in most decoding steps. These results further demonstrate the effectiveness of LogitSpec.
[…]
Method Spec-Bench HumanEval
Vicuna 7B Vicuna 13B Vicuna 33B Vicuna 7B Vicuna 13B Vicuna 33B
PLD 63.88 63.08 62.71 69.03 69.51 67.67
LogitSpec 97.76↑53.04% 97.64↑54.79% 97.93↑56.16% 99.29↑43.84% 99.31↑42.87% 99.37↑46.84%
6 Conclusion
Conclusion.
[…]
Limitations and ethic statements.
While our LogitSpec is a fully plug-and-play SD framework, it is possible to integrate a draft model for further acceleration. We affirm our commitment to contributing positively to society, avoiding harm, and upholding honesty and trustworthiness. We appropriately cite the previous methods and datasets we use, and ensure that all data involved is fully public, with no private data being utilized.
Appendix A More Discussions to Related Work
We give more discussions to existing works in both draft-model-based speculative decoding and draft-model-free speculative decoding.
Draft-model-based speculative decoding.
Besides the discussions in Section 2, there are also many excellent works related to the draft-model-based speculative decoding. For example, HASS [Zhang et al., 2025] discovers the inconsistency of EAGLE between training and inference, and proposes a multi-step training framework to address this. CORAL [Weng et al., 2025] proposes a cross-step representation alignment to address this problem. Judge Decoding [Bachmann et al., 2025] recognizes the potential of accepting high-quality but refused draft tokens to further improve the acceleration. Gumiho [Li et al., 2025c] demonstrates that the initial draft token is more important and proposes a hybrid model to combine serial and parallel draft head.
Besides these methods that focus on the training process, the process of verification also draws extensive interests, mainly focus on adaptive draft length. SpecDec++ [Huang et al., 2025] formulates the verification process as a Markov decision process to adaptively determine the draft length. OPT-Tree [Wang et al., 2024] proposes a method to search the optimal tree structure that maximizes the mathematical expectation of the acceptance length in each decoding step. AdaEAGLE [Zhang et al., 2024b] proposes a novel framework to explicitly model the draft tree structure for EAGLE series models. PEARL [Liu et al., 2025] pioneers this direction by serving the draft model and the target model in parallel to achieve a segmented adaptive draft length.
Draft-model-free speculative decoding.
For layer sparsity, SPEED [Hooper et al., 2024] proposes a method to speculatively execute multiple future tokens in parallel with the current token using predicted values based on early-layer hidden states. FREE [Bae et al., 2023] proposes a shallow-deep module and a synchronized parallel decoding to improve the efficiency. EESD [Liu et al., 2024] proposes an early-exiting framework with a self-distillation method and leverages Thompson Sampling to regulate the generation processes. For retrieval-based SD, Token Recycling [Luo et al., 2024] proposes a method to utilize the dropped draft tokens and generate draft tokens via an adjacency matrix. Different with our method, Token Recycling is a query-dependent method that utilize information from other queries, which may result in limitations when applied to real-world applications with complex and dynamic user inputs. SAM Decoding [Hu et al., 2024] utilizes common text corpus and dynamic text sequence as retrieved sources and proposes suffix automaton to efficiently obtain yields more accurate match positions.
Appendix B Experimental Setup for the Motivation Experiments
To investigate the prediction ability of the last logit, we conduct two motivation experiments to demonstrate the effectiveness and robustness of the last logit. For the effectiveness of the last logit, as shown in Figure 2, we conduct auto-regressive inference for the 6 models and record the logits for each decoded token. Then, for each decoded token $x_{i}$ which is sampled from the logit $p_{i-1}$, we investigate the rank of $x_{i}$ in $p_{i-2}$, i.e., the corresponding last logit, and visualize the statistics of the rank in Figure 2. For the robustness of the last logit, as shown in Figure 2, we conduct last logit decoding and investigate its mean accepted tokens per decoding step. Both the two experiments are conducted on a small subset of Spec-Bench, where we randomly sample 2 questions for each sub-category of MT, and 10 questions for other categories (Trans, Sum, QA, Math, RAG).
Appendix C Pseudo Code to Prepare Attention Inputs
We give a pseudo code to organize the retrieval multiple draft token sequences into a draft tree and prepare its attention mask. More detailed implementations can be found in our attached code.
def prepare_attention_inputs(past_len, next_token, candidate_list, num_draft_tokens):
’LogitSpec organizes draft tokens in a tree manner. Each sub-sequence corresponds to a local causal mask.’
seq_len = num_draft_tokens + 1
# organize the candidate list into a sequence
draft_ids = [next_token] + [token for sub in candidate_list for token in sub]
# prepare original position ids and attention mask
position_ids = torch.zeros((1, seq_len), dtype=torch.long)
causal_mask = torch.full((seq_len, past_len + seq_len), fill_value=0)
causal_mask[:, :past_len+1] = 1
# prepare causal mask
idx = 1
for sub_sequence in candidate_list:
l = len(sub_sequence)
sub_mask = torch.tril(torch.ones((l, l)))
causal_mask[idx:idx+l, idx+past_len:idx+past_len+l] = sub_mask
position_ids[0, idx:idx+l] = torch.arange(l) + 1
idx += l
position_ids += past_len
return draft_ids, causal_mask, position_ids
Appendix D More Implementation Details
D.1 Retrieval Process
The retrieval process of retrieval-based SD methods will affect the overall acceleration. While the existing method PLD retrievals reference with string matching, it is unaffordable for LogitSpec to retrieval in this way. Specifically, suppose the matching ngram size is $m$, the length of the whole token sequence is $n$, string matching takes the time complexity of $O(mn)$. However, as LogitSpec needs to retrieval for all the next next tokens, the retrieval takes the time complexity of $O(kmn)$, where $k$ is the number of the next next tokens.
To address this issue, we propose a simple yet effective method to reduce the retrieval overhead by constructing a hash table. Specifically, we go through the token sequence, storing each key ngram with size $1$ to $m$ and its value ngram (its following tokens). In this way, the retrieval overhead is reduced to $O(n+k)$ with an additional memory cost $O(mn)$. As the length of the token sequence is relatively small, the additional memory cost is negligible. Furthermore, we will update the retrieval model $\mathcal{R}$ with the decoded tokens in the same way during each decoding step.
D.2 Evaluation Instructions
In our experiments, we employ different instructions for different evaluation tasks. Specifically, for Spec-Bench, we use its standard instructions:
Prompt Templates for Spec-Bench A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: Question
ASSISTANT:
For CNN/DM, we add a simple instruction “Summarize: ” to the instruction:
Prompt Templates for CNN/DM A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: Summarize: Question
ASSISTANT:
For GSM8K, we follow the setting with [Liu et al., 2025] and use a 8-shot CoT for inference:
Prompt Templates for GSM8K A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: 8-shot CoT Q: Question
ASSISTANT:
For HumanEval, we add a simple instruction “Please help me to complete this code, just output your codes directly.”:
Prompt Templates for HumanEval A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: Please help me to complete this code, just output your codes directly. Question
ASSISTANT:
D.3 Dataset Configurations
In our experiments, we evaluate the effectiveness of our LogitSpec on 4 categories of text generation tasks, including Spec-Bench, CNN/DM, GSM8K, and HumanEval. For Spec-Bench and HumanEval, we use the full data for evaluation. For CNN/DM and GSM8K, we randomly sample 1000 questions for evaluation. The maximal generation length is set as 1024 across all the experiments.
D.4 Model Configurations
In our experiments, all the models are deployed with precision float16. The <|eos token|> is set the same as the tokenizer’s <|eos token|>. To effectively alleviate the overhead of rollbacking KV cache to the accepted draft tokens, we follow Medusa [Cai et al., 2024] to allocate continuous GPU memory for all the KV cache.
Appendix E More Experiment Results
Here, we present the results of Llama 2 chat series models on Spec-Bench, CNN/DM, GSM8K, and HumanEval in Table 5 and Table 6.
Table 5: Experimental results of LogitSpec on CNN/DM [Nallapati et al., 2016], GSM8K [Cobbe et al., 2021] and HumanEval [Chen et al., 2021] with Llama-2. We report the mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
Models Method CNN/DM GSM8K HumanEval Overall
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Llama 2 7B Lookahead 1.58 1.36 2.02 1.67 1.77 1.53 1.79 1.52
REST 1.71 1.33 1.51 1.18 1.97 1.52 1.73 1.34
PLD 1.89 1.73 3.32 2.98 1.57 1.41 2.26 2.04
SpS 1.99 1.46 2.83 1.87 2.13 1.51 2.32 1.61
LogitSpec 2.41 2.02 4.44 3.68 2.17 1.75 3.01 2.48
Llama 2 13B Lookahead 1.56 1.18 2.08 1.52 1.84 1.66 1.83 1.45
REST 1.71 1.34 1.53 1.26 1.96 1.63 1.73 1.41
PLD 1.89 1.52 3.24 2.54 1.73 1.63 2.29 1.90
SpS 1.95 1.34 2.87 1.85 2.33 1.76 2.38 1.65
LogitSpec 2.43 2.03 4.31 3.24 2.38 2.10 3.04 2.46
Llama 2 70B Lookahead 1.53 1.28 1.90 1.57 1.86 1.57 1.76 1.47
REST 1.67 1.35 1.63 1.32 1.96 1.66 1.75 1.44
PLD 1.98 1.74 1.63 1.46 1.62 1.49 1.74 1.56
SpS 2.01 1.71 1.98 1.69 2.21 1.70 2.07 1.70
LogitSpec 2.67 2.10 2.37 1.87 2.33 1.93 2.46 1.97
Table 6: Experimental results of LogitSpec on Spec-Bench [Xia et al., 2024] with Llama-2. We report the speedup ratio on each sub task, mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
Models Method MT Trans Sum QA Math RAG MAT Speedup
Llama 2 7B Lookahead 1.55 1.44 1.42 1.48 1.69 1.46 1.69 1.51
REST 1.58 1.20 1.38 1.61 1.31 1.55 1.83 1.48
PLD 1.46 1.34 1.89 1.23 1.65 1.58 1.59 1.49
SpS 1.57 1.38 1.55 1.46 1.55 1.60 2.07 1.53
LogitSpec 1.83 1.72 2.19 1.63 2.15 1.94 2.15 1.87
Llama 2 13B Lookahead 1.39 1.36 1.21 1.39 1.69 1.25 1.68 1.37
REST 1.53 1.14 1.30 1.51 1.23 1.45 1.85 1.42
PLD 1.36 1.19 1.58 1.16 1.62 1.37 1.56 1.37
SpS 1.52 1.26 1.43 1.42 1.54 1.50 2.03 1.48
LogitSpec 1.73 1.57 1.86 1.53 2.14 1.73 2.12 1.74
Llama 2 70B Lookahead 1.45 1.35 1.28 1.38 1.71 1.31 1.66 1.41
REST 1.63 1.33 1.38 1.67 1.35 1.55 1.83 1.53
PLD 1.34 1.32 1.76 1.18 1.63 1.47 1.51 1.39
SpS 1.65 1.50 1.62 1.57 1.70 1.68 1.88 1.63
LogitSpec 1.66 1.58 1.95 1.58 2.03 1.78 2.12 1.72
Appendix F Broader Impact
Social Impact. This paper investigates a retrieval-based speculative decoding technique for LLM inference acceleration, enabling more discussions and potential efficient discovery of novel methods. Our LogitSpec enhances LLMs’ efficiency on downstream tasks and potentially improves user’s satisfaction on LLM-based applications.