### v2

# LogitSpec: Accelerating Retrieval-based Speculative Decoding via Next Next Token Speculation

Abstract
[…]
LogitSpec: Accelerating Retrieval-based Speculative Decoding via Next Next Token Speculation
Tianyu Liu1,2thanks: Equal Contribution.   Qitan Lv1,2∗   Hao Li1,2   Xing Gao2   Xiao Sun2thanks: The Corresponding Author.   Xiaoyan Sun1 1University of Science and Technology of China  2Shanghai AI Laboratory {tianyu_liu, qitanlv, lihaohn}@mail.ustc.edu.cn sunxiao@pjlab.org.cnsunxiaoyan@ustc.edu.cn
1 Introduction
Large Language Models (LLMs), such as GPT-4.5 (OpenAI, 2024), DeepSeek R1 (DeepSeek-AI, 2025), Qwen 2.5 (Team, 2025), and LLaMA-3 (Team, 2024), have demonstrated remarkable capabilities across a wide range of natural language processing tasks, including question answering (Calijorne Soares and Parreiras, 2020), code generation (Jiang et al., 2024), and dialogue systems (Yi et al., 2024). While they achieve success by scaling up the model parameters, the increase in scale comes with significant inference challenges. The most straightforward challenge is auto-regressive decoding, where each LLM decoding step can only generate one token. This token-by-token decoding process incurs exacerbated latency with both the length of generated tokens and the model scale.
To address this challenge, speculative decoding (SD) (Leviathan et al., 2023; Chen et al., 2023) has emerged as a promising approach for lossless LLM inference acceleration. The key idea of SD is to employ a small draft model to first generate $\gamma$ draft tokens, and then the target model validates these tokens in parallel within a single forward, ensuring the output distribution to be unchanged. However, deploying an extra draft model introduces considerable overhead and difficulties, including (a) complex implementation and deployment, especially in integration with other techniques,(Liu et al., 2025) (b) additional memory overhead, especially in long-context scenarios. Moreover, when there exists no available draft model for SD, it takes substantial training cost to construct a compact draft model (Li et al., 2025b). (Please refer to Section 2 for more details.)
Therefore, many existing works are dedicated to developing draft-model-free SD methods (Saxena, 2023; Fu et al., 2024; He et al., 2024), where the draft tokens are generated in a retrieval-based manner. At each decoding step, they retrieve the relevant reference according to the last few tokens and extract draft tokens from the reference. In this way, they can explicitly eliminate the need for a draft model and reduce the drafting overhead.
Figure 1: Illustration of vanilla retrieval-based SD method and our LogitSpec. (a) Retrieval-based SD retrieves the wrong token “triangle” according to the next token “the”. (b) LogitSpec first speculates the next next token “area”, and then retrieves the right relevant reference “area of” according to “the area”. This simple example illustrates how LogitSpec utilizes the last logit to speculate the next next token and improves the retrieval accuracy.
Despite their promising efficacy and implementation simplicity, these methods rely on a matching paradigm to effectively retrieve the most relevant reference as draft tokens and often struggle to find matched and accurate tokens. As shown in Figure 1 (a), with a simple prompt “What is the area of the triangle?”, vanilla retrieval-based SD methods fail to effectively retrieve the right token “area” according to the next token “the”, as the most relevant reference is “the triangle”. Moreover, it is often the case that no matched reference can be found, e.g., there are no matched tokens in more than 30% of decoding steps in PLD (Saxena, 2023).
In this paper, we propose a simple yet effective retrieval-based SD framework, namely LogitSpec, which leverages the logit of the last token (last logit) to predict the next next token and improves the accuracy of the retrieval reference. LogitSpec is motivated by the observation that the last logit can speculate the next next token with a relatively high accuracy. Based on this observation, we use the speculated next next token as guidance to retrieve reference. Specifically, LogitSpec generates draft tokens in two steps: (1) utilizing the last logit to speculate the next next token; (2) retrieving relevant reference for both the next token and the next next token. As shown in Figure 1 (b), with the speculated token “area”, LogitSpec successfully retrieves the correct draft tokens “area of”. This two-step process enables better retrieval accuracy—it can help filter the most relevant reference when the reference is redundant, while extending the searching space when there exists no relevant reference. To summarize, our contributions are:
  • (a)
    We empirically observe that the last logit can speculate the next next token with a relatively high accuracy. Unlike other retrieval techniques, which are sensitive to the specific task, this property is robust and effective across various tasks.
  • (b)
    Building on this observation, we propose LogitSpec, a plug-and-play retrieval-based SD framework which can improve the retrieval accuracy and achieve better speedup.
  • (c)
[…]
2 Related Work
In this section, we briefly review the two main directions of speculative decoding, draft-model-based speculative decoding and draft-model-free speculative decoding. More discussions are provided in Appendix A.
Draft-model-based speculative decoding.
This line of research focuses on improving the draft quality, which primarily invests heavily in post-training a specialized draft model to generate high-quality drafts. Medusa (Cai et al., 2024) pioneers this direction by adding extra MLP heads at the top of the target model, reusing the hidden states from the target model, and predicting the next few tokens in parallel. However, the generation from these decoding heads is independent, which harms the accuracy of the draft tokens. Therefore, some works (Ankner et al., 2024; Xiao et al., 2024) propose to add sequential dependencies for better performance. Glide (Du et al., 2024) takes a similar idea of Medusa and proposes to reuse the KV cache from the target model. Recently, Eagle (Li et al., 2025c, 2024, b) has dominated this research line by training an auto-regressive head and scaling up the training data. However, while these methods achieve superior speedup, all of them rely on an extra draft model, which necessitates extra parameters or extensive training. The existence of the draft model significantly limits its application: (a) deploying a draft model requires complex and careful implementation. (b) Both the draft model and its KV cache require additional GPU memory, which may incur load imbalance. In the long context settings, even the KV cache for the draft model may exceed the capacity of 1 GPU and lead to significant resource competitions; (c) the training cost is substantial, and the draft model requires retraining when the target model is updated.
Draft-model-free speculative decoding.
This line of research focuses on maximizing drafting efficiency by eliminating the need for a draft model entirely, aiming for universal, zero-cost acceleration. Some existing works (Zhang et al., 2024a; Elhoushi et al., 2024; Xia et al., 2025) observe that the layer sparsity of LLMs means that not all the layers are necessary to predict the next token. They propose methods to generate draft tokens with a subset of the layers of the target model. However, the number of skipped layers is limited, which affects the overall speedup. Recently, retrieval-based SD (He et al., 2024; Saxena, 2023; Fu et al., 2024) has dominated the draft-model-free speculative decoding by retrieving the relevant reference as the draft tokens. The relevant reference can be derived from the question and generated tokens, or an external database. However, there exist two common issues during the retrieval process: (a) retrieval from the context often fails to find matched draft tokens; (b) retrieval from an external database often struggles to find accurate draft tokens. These two issues motivate us to develop a more appropriate retrieval-based speculative decoding framework.
3 Background
Speculative Decoding.
Let $\bm{x}$ denote an input sequence (prefix). A speculative decoding step consists of a drafting phase and a verification phase. During the drafting phase, the draft model $\mathcal{M}_{q}$ is employed to give $\gamma$ draft tokens $x_{1},x_{2},...,x_{\gamma}$ by running $\gamma$ times the model forward and sampling. Here, we denote the output logit $\mathcal{M}_{q}(\bm{x}+[x_{1},...,x_{i-1}])$ as $q_{i-1}$, then each draft token is given by $x_{i}\sim q_{i-1},i=1,...,\gamma$. During the verification phase, the prefix $\bm{x}$ together with $\gamma$ draft tokens are sent to $\mathcal{M}_{p}$ for verification. The target model $\mathcal{M}_{p}$ inputs $\bm{x}+[x_{1},...,x_{\gamma}]$ and outputs the logits $p_{0},p_{1},...,p_{\gamma}$. Then SD sequentially verifies $x_{i}$ via speculative sampling (Leviathan et al., 2023; Chen et al., 2023), where the acceptance rate is given by:
[…]
If SD rejects $x_{i}$, it will resample a token from $norm(\max(0,p_{i-1}-q_{i-1}))$, otherwise, SD accepts all the draft tokens and samples an additional token from $p_{\gamma}$. In this way, each SD step generates at least 1 token and at most $\gamma+1$, leading to theoretical lossless quality and efficiency acceleration.
Retrieval-based Speculative Decoding.
The retrieval-based SD methods generate draft tokens in a retrieval-based manner. As retrieval-based SD methods do not require an additional draft model, we omit the notation for $\mathcal{M}_{q}$ and abbreviate $\mathcal{M}_{p}$ as $\mathcal{M}$. A retrieval model $\mathcal{R}$ is employed to store $n$-grams (referred as reference) from corpus
$\mathcal{R}=\{(x_{i}^{1},x_{i}^{2},...,x_{i}^{n})\}_{i=1}^{N}.$ (2)
Here, $N$ denotes the total number of $n$-grams in $\mathcal{R}$. Existing retrieval-based SD methods mainly differ in constructing $\mathcal{R}$. At each SD decoding step, given a query $m$-gram tokens $x_{q}^{1},x_{q}^{2},...,x_{q}^{m}$, the retrieval model first traverses the reference, finds matched $n$-grams and returns the subsequent tokens of the matched tokens:
$\textsc{Match}\big(\mathcal{R},(x_{q}^{1},x_{q}^{2},...,x_{q}^{m})\big)\\ =\{(x_{i}^{m+1},...,x_{i}^{n})\mid i\in\mathcal{S}\}.$ (3)
where $\mathcal{S}=\{i|x_{i}^{t}=x_{q}^{t},\forall t=1,...,m\}$. Intuitively, a larger $m$ leads to more precise matches but lower match probabilities.
4 Method
Figure 2: Motivated observation I: the last logit can speculate the next next token with high accuracy. In over 50% decoding steps, the next next token can be found in the top-60 entries within the last logit across model sizes and architectures.
In this section, we introduce our LogitSpec, a novel retrieval-based SD framework that utilizes last logit as a guidance for retrieval and effectively improves the retrieval performance. We first conduct a motivated experiment in Section 4.1, and then introduce the whole framework of LogitSpec in Section 4.2 and 4.3. An overview of LogitSpec is shown in Figure 4.
4.1 Motivated Observation
Drawing insights from multi-token prediction (Gloeckle et al., 2024) that fine-tuned LLMs can predict multiple tokens in a single forward, we are interested in the ability of LLMs to predict the next next token without fine-tuning. We conduct a simple experiment to investigate the rank of the next next token in the last logit, which is used to predict the next token. Specifically, we use a small fraction of Spec-Bench (Xia et al., 2024) that contains 13 sub-tasks (detailed in Appendix B) and 6 different LLMs as backbones. Results in Figure 2 demonstrate that the last logit has a strong potential to predict the next next token. In over 50% decoding steps, the next next token can be found within the top-60 entries of the last logit. For Llama 3.1 8B (Team, 2024) with a large 128K vocabulary, this percentage even increases to 64% within the top-60 entries of the last logit. The results of 6 different LLMs on 13 downstream tasks demonstrate that this observation holds consistently across different models (Vicuna, Llama 3.1, and Qwen) and scales (ranging from 0.5B to 33B).
Figure 3: Motivated observation II: compared with other retrieval-based methods, the prediction of the last logit demonstrates robustness to downstream tasks, motivating us to utilize it to guide the retrieval process.
We further conduct experiments on Spec-Bench with Vicuna 7B to investigate the robustness of the predictive capability of the last logit across different tasks. We propose a minimal speculative decoding algorithm with last logit, last logit decoding for this experiment, where we directly use the top-60 entries of the last logit to serve as the draft tokens. Each draft token is a guess at the next next token. If the target model accepts any one of the draft tokens, the others are dropped. Consequently, at each decoding step, the accepted length is either 2 or 1. As in Figure 3, this simple SD method yields mean accepted tokens per decoding step (MAT) of over 1.5. Compared with other retrieval-based SD methods such as PLD (Saxena, 2023) and REST (He et al., 2024), which are highly affected by the task type, the MAT of last logit decoding exhibits superior robustness to the task. These results demonstrate that the ability of the last logit to predict the next next token widely exists in modern LLMs themselves, and strongly motivate us to improve retrieval-based SD methods with the last logit. We provide a detailed experimental setup for the motivation experiments in Appendix B.
Figure 4: An overview of LogitSpec. At each decoding step, LogitSpec first utilizes the top-$k$ entries of the last logit as the speculation for the next next token. Then, LogitSpec retrieves relevant references for both the next token and the next next token. Finally, LogitSpec organizes the draft tokens into a draft tree and prepares a tree attention for parallel verification.
4.2 LogitSpec Drafting
While last logit decoding has shown great simplicity in implementation and robustness to the downstream tasks, its theoretical upper bound of the speedup ratio is 2, as all the draft tokens are guesses for the next next token. In our experiments, the last logit decoding achieves an overall speedup ratio of $1.2\times\sim 1.4\times$, which hinders the real-world applications of last logit decoding due to the unsatisfactory speedup ratio. Therefore, to further improve the overall inference speedup, we propose to utilize the prediction ability of the last logit as a guidance for retrieval. Specifically, for a given input prefix $\mathbf{x}=(x_{1},x_{2},...,x_{i})$ and the last logit $p_{i}=\mathcal{M}(x_{1},x_{2},...,x_{i})$, we first sample the next token $x_{i+1}\sim p_{i}$. Then, we utilize the top-$k$ entries of the last logit (except the next token $x_{i+1}$) as the speculation to the next next token $\widetilde{x}_{i+2}$:
$\mathcal{L}=\{\widetilde{x}_{i+2}\mid\widetilde{x}_{i+2}\in\textsc{Top}_{k}(p_{i}),\,\widetilde{x}_{i+2}\neq x_{i+1}\}.$ (4)
Then, we use the retrieval model $\mathcal{R}$ to retrieve reference for both the next token and next next tokens:
$\mathcal{D}_{i+1}(x_{i+1})=\textsc{Match}\big(\mathcal{R},\bm{g}^{(1)}_{i}(x_{i+1})\big),$ (5)
$\mathcal{D}_{i+2}(\widetilde{x}_{i+2})=\textsc{Match}\big(\mathcal{R},\bm{g}^{(2)}_{i}(\widetilde{x}_{i+2})\big),$ (6)
$\mathcal{D}=\mathcal{D}_{i+1}(x_{i+1})\cup\bigcup_{\widetilde{x}_{i+2}\in\mathcal{L}}\mathcal{D}_{i+2}(\widetilde{x}_{i+2}).$ (7)
Here, $\bm{g}^{(1)}_{i}(x_{i+1})=(x_{i-m+2},...,x_{i},x_{i+1})$ denotes the $m$-gram query for the next token, and $\bm{g}^{(2)}_{i}(\widetilde{x}_{i+2})=(x_{i-m+3},...,x_{i},x_{i+1},\widetilde{x}_{i+2})$ denotes the query for the next next token. In our experiments, we retrieve draft tokens with $m=3$ first. If no matched tokens are found, we decrease $m$ to 2 and so on. For simplicity, we only use the user-input prompt and decoded tokens to construct the retrieval model $\mathcal{R}$, that is to say, our LogitSpec is query-independent, and will not be affected by other queries. We give the detailed implementation of the retrieval process in Appendix D.3.
Taking the example in Figure 4, we first select the top-$k$ entries of the last logit as the speculation for the next next token. Then, we retrieve reference for both the next token “the” and the next next token “area”, and return “of the”. After that, we organize these draft tokens into a draft tree, where the next token serves as the root. Finally, the token sequence “area of the” is accepted by $\mathcal{M}$.
After the retrieval process, we employ a simple pruning strategy to control the number of draft tokens. Specifically, we do not prune the retrieval tokens for the next token. For each speculated next next token, we prune the retrieval tokens with a simple heuristic strategy: if the rank of the next next token is below 8, we preserve 4 tokens; if the rank of the next next token is below 32, we preserve 3 tokens; otherwise, we only preserve the speculated next next token itself. The retrieval process is terminated until the total number of draft tokens exceeds a specific draft tree capacity K{K}. Furthermore, we explored the impact of various pruning strategies, with results detailed in Table 14 of Appendix E.7.
4.3 LogitSpec Verification
After retrieving multiple draft token sequences $\mathcal{D}$, we organize these token sequences into a draft tree and prepare a tree attention for parallel verification.
Then, we prepare an attention mask to make each draft token sequence invisible to other sequences. Let $\Lambda(l)$ denote a causal mask with length $l$, and $m_{j}$ denote the length of the $j$-th token sequence in $\mathcal{D}$; the attention mask is given by:
$A_{\text{draft}}=\operatorname{diag}\big(\Lambda(m_{1}),\ldots,\Lambda(m_{j})\big).$ (8)
An illustration of this attention mask is shown in Figure 4(c) (note that the next token is visible to all sequences, corresponding to the first column). Both the preparation of the attention mask and the verification of the draft tree are consistent with previous works (Miao et al., 2024; Cai et al., 2024). We provide a pseudo code of the attention mask preparations in Appendix C.
5 Experiments
5.1 Experimental Setup
Tasks and Datasets.
We evaluate LogitSpec on a broad suite of text generation benchmarks. We first use Spec-Bench (Xia et al., 2024), a widely used comprehensive benchmark that covers diverse application scenarios, including multi-turn conversation (MT), translation (Trans), summarization (Sum), question answering (QA), mathematical reasoning (Math), and retrieval-augmented generation (RAG). We then further evaluate LogitSpec on HumanEval (Chen et al., 2021), GSM8K (Cobbe et al., 2021), CNN/DM (Nallapati et al., 2016), MATH (Hendrycks et al., 2021), AIME 24&25 (Mathematical Association of America, 2025), and LongBench (Bai et al., 2024), which are widely used benchmarks for code generation, math reasoning, summarization, and long-context generation.
Implementation Details.
We follow standard inference settings with temperature 0 and batch size 1. All implementation details (baselines, hardware, software versions, model configurations) are provided in Appendix D. We use two widely used metrics for evaluation: mean accepted tokens per decoding step (denoted as MAT) and overall speedup ratio (denoted as Speedup). We further evaluate LogitSpec under non-zero temperatures ($T>0$) in Appendix E.4.
5.2 Main Results
We conduct experiments on various text generation benchmarks to demonstrate the effectiveness of LogitSpec. As in Table 1 (with the Spec-Bench breakdown deferred to Appendix E.2), LogitSpec outperforms retrieval-based SD baselines by a large margin. We further present more in-depth analysis of LogitSpec including (i) more LLM backbones (Llama2 series, Llama-3.1-Instruct-8B and Qwen-3-8B), (ii) more benchmarks (MATH and AIME datasets), and (iii) long context benchmarks (the LongBench dataset) in Appendix E.
Table 1: Experimental results of LogitSpec on CNN/DM (Nallapati et al., 2016), GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) with Vicuna. We report the mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
Models Method CNN/DM GSM8K HumanEval Spec-Bench Average
MAT Speedup MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Vicuna 7B Lookahead 1.54 1.28 1.92 1.64 1.79 1.57 1.66 1.27 1.73 1.44
REST 1.64 1.19 1.55 1.13 1.96 1.51 1.82 1.48 1.74 1.33
PLD 2.61 2.26 1.80 1.62 1.84 1.65 1.73 1.59 2.00 1.78
SpS 2.34 1.58 2.08 1.49 2.56 1.80 2.28 1.49 2.32 1.59
LogitSpec 3.28 2.61 2.69 2.20 2.62 2.24 2.44 2.01 2.76 2.26
Vicuna 13B Lookahead 1.46 1.12 1.88 1.61 1.75 1.57 1.63 1.22 1.68 1.38
REST 1.65 1.22 1.57 1.18 1.94 1.55 1.82 1.38 1.75 1.33
PLD 2.34 1.85 1.76 1.61 1.98 1.77 1.68 1.48 1.94 1.68
SpS 2.18 1.48 2.00 1.55 2.66 1.95 2.18 1.49 2.26 1.62
LogitSpec 2.90 2.17 2.59 2.13 2.88 2.47 2.32 1.93 2.67 2.18
Vicuna 33B Lookahead 1.50 1.18 1.89 1.44 1.66 1.33 1.61 1.24 1.67 1.30
REST 1.65 1.41 1.57 1.32 1.99 1.64 1.80 1.48 1.75 1.46
PLD 2.07 1.72 1.66 1.41 1.60 1.42 1.55 1.42 1.72 1.49
SpS 2.01 1.57 1.97 1.54 2.27 1.80 2.01 1.61 2.07 1.63
LogitSpec 2.66 2.01 2.46 1.93 2.41 1.95 2.13 1.76 2.42 1.92
[…]
5.3 Ablation Study
To further provide more insights of LogitSpec, we conduct extensive ablation studies on the aforementioned 4 benchmarks in Table 2. Specifically, we mainly focus on two components of LogitSpec: last logit and retrieval. We denote LogitSpec without the last logit decoding and only using retrieval as w/o last logit, and LogitSpec without retrieval and only using last logit decoding as w/o retrieval. As shown in Table 2, the absence of any component results in a performance degradation of the entire framework.
Our findings are as follows. First, the absence of retrieval exhibits more importance to the final acceleration, which is consistent with the discussion in Section 4.2 that the theoretical upper bound of last logit decoding severely hinders its real-world application. Second, the absence of retrieval and the absence of last logit show different effects in different sub-tasks. For example, the absence of retrieval decreases MAT by 0.87 on Spec-Bench with Vicuna 7B, while it decreases MAT by 1.71 on CNN/DM. In contrast, the absence of last logit leads to a more consistent MAT degradation across different tasks, which further highlights the robustness of last logit decoding. Finally, these results demonstrate that combining last logit with retrieval improves the retrieval performance and overall speedup. We also conduct ablation studies on pruning strategies in Appendix E.7 and E.8. The experimental results suggest that varying the pruning strategy yields only minor differences.
[…]
Models Method Spec-Bench CNN/DM GSM8K HumanEval
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Vicuna 7B w/o last logit 1.72↓.72 1.27↓.74 2.66↓.62 2.24↓.37 1.81↓.88 1.63↓.57 1.89↓.73 1.69↓.55
w/o retrieval 1.57↓.87 1.24↓.77 1.57↓1.71 1.23↓1.38 1.71↓.98 1.39↓.81 1.69↓.93 1.39↓.85
LogitSpec 2.44 2.01 3.28 2.61 2.69 2.20 2.62 2.24
Vicuna 13B w/o last logit 1.66↓.66 1.23↓.70 2.34↓.56 1.81↓.36 1.78↓.81 1.61↓.52 2.02↓.86 1.75↓.72
w/o retrieval 1.59↓.73 1.19↓.74 1.58↓1.32 1.22↓.95 1.71↓.88 1.45↓.68 1.63↓1.25 1.37↓1.10
LogitSpec 2.32 1.93 2.90 2.17 2.59 2.13 2.88 2.47
Vicuna 33B w/o last logit 1.55↓.58 1.22↓.54 2.05↓.61 1.72↓.29 1.64↓.82 1.39↓.54 1.59↓.82 1.40↓.55
w/o retrieval 1.61↓.52 1.25↓.51 1.62↓1.04 1.26↓.75 1.69↓.77 1.33↓.60 1.60↓.81 1.34↓.61
LogitSpec 2.13 1.76 2.66 2.01 2.46 1.93 2.41 1.95
[…]
Method Spec-Bench HumanEval
Vicuna 7B Vicuna 13B Vicuna 33B Vicuna 7B Vicuna 13B Vicuna 33B
PLD 63.88 63.08 62.71 69.03 69.51 67.67
LogitSpec 97.76↑53.04% 97.64↑54.79% 97.93↑56.16% 99.29↑43.84% 99.31↑42.87% 99.37↑46.84%
5.4 Case Study
Figure 5: Running time breakdown of the whole decoding process on Spec-Bench with Vicuna 7B.
In-depth Running Time Analysis.
We conduct experiments to analyze the running time allocation of the whole decoding. Specifically, there are five non-negligible components in LogitSpec, including (a) retrieving draft tokens: the process of retrieving reference for the next token and the next next tokens; (b) preparation: preparing attention mask for the draft tokens; (c) model forward: conducting one-pass model forward; (d) verification: validating the draft tokens with speculative sampling; (e) update: necessary update of KV cache and retrieval model.
As shown in Figure 5, model forward occupies the majority of wall-clock time. Compared with vanilla AR decoding, the overhead introduced by LogitSpec, i.e., retrieving overhead, only takes 1.17% of the whole decoding process. It can be further alleviated by parallel techniques, as the retrieval process is independent, which brings negligible overhead as well.
Retrieval Performance.
As mentioned in Section 4.2, LogitSpec expands the retrieval range and improves the retrieval performance. We conduct a simple experiment to test the retrieval success rate, i.e., whether the model successfully retrieves the reference as draft tokens. As in Table 3, PLD cannot retrieve any matched tokens in more than 30% decoding steps, while LogitSpec retrieves matched reference in most decoding steps. These results further demonstrate the effectiveness of LogitSpec.
Real-world examples.
We also provide a real-world example in Appendix E.9 to illustrate how next-next token speculation allows LogitSpec to succeed where standard retrieval methods fail.
6 Conclusion
In this paper, we observe that the logit of the last token can predict the next next token with a relatively high accuracy without any fine-tuning. Based upon this observation, we propose a novel retrieval-based SD framework, namely LogitSpec, which utilizes the prediction ability of the last logit to effectively expand the retrieval range and find the most relevant reference as the draft tokens. LogitSpec does not require an additional draft model and is a fully plug-and-play method, which can be easily implemented and integrated into existing LLM frameworks. Extensive experiments demonstrate that LogitSpec can effectively improve the retrieval performance, leading to a $1.8\times\sim 2.6\times$ speedup across all the evaluation benchmarks.
7 Limitations and Future Work.
We consider a few limitations and future works as follows. (i) While our LogitSpec is a fully plug-and-play SD framework, its real-world inference acceleration is less competitive. Our future works involve integrating LogitSpec into existing draft-model-based SD methods for further acceleration. (ii) Currently, LogitSpec retrieves relevant reference from the prompt, which may incur lower speedup when the prompt is short. We consider integrating an external database to boost the retrieval model as a future work. (iii) Our acceleration estimates rely on offline decoding with curated prompts and NVIDIA A100 GPUs, so the reported speedups may vary on other workloads or hardware. LogitSpec currently retrieves draft candidates from the on-the-fly prompt context only; short prompts or out-of-domain corpora reduce its benefits.
8 Ethic Statements
Inheritance of Model Behaviors.
LogitSpec is an inference acceleration framework designed to speed up existing Large Language Models without modifying their weights. As a speculative decoding method, it aims to losslessly recover the distribution of the target model. Consequently, LogitSpec inherits the ethical properties, biases, and potential safety risks of the underlying target LLM and draft model. It does not introduce new capabilities for generating harmful content, nor does it mitigate existing biases in the base models. Users should continue to apply standard safety guardrails and alignment techniques to the target models deployed with LogitSpec.
Appendix A More Discussions to Related Work
We provide additional discussion of existing works in both draft-model-based speculative decoding and draft-model-free speculative decoding. An intuitive comparison is shown in Table 4.
Draft-model-based speculative decoding.
Besides the discussions in Section 2, we discuss additional works on draft-model-based speculative decoding. For example, HASS (Zhang et al., 2025) identifies the inconsistency between EAGLE’s training and inference phases, and proposes a multi-step training framework to address this. CORAL (Weng et al., 2025) proposes a cross-step representation alignment to address this problem. Judge Decoding (Bachmann et al., 2025) recognizes the potential of accepting high-quality but refused draft tokens to further improve the acceleration. Gumiho (Li et al., 2025a) demonstrates that the initial draft token is more important and proposes a hybrid model to combine serial and parallel draft heads.
Besides these methods that focus on the training process, the process of verification also draws extensive interest, mainly focusing on adaptive draft length. SpecDec++ (Huang et al., 2025) formulates the verification process as a Markov decision process to adaptively determine the draft length. OPT-Tree (Wang et al., 2024) proposes a method to search for the optimal tree structure that maximizes the mathematical expectation of the acceptance length in each decoding step. AdaEAGLE (Zhang et al., 2024b) proposes a novel framework to explicitly model the draft tree structure for EAGLE series models. PEARL (Liu et al., 2025) pioneers this direction by serving the draft model and the target model in parallel to achieve a segmented adaptive draft length.
Draft-model-free speculative decoding.
For layer sparsity, SPEED (Hooper et al., 2024) proposes a method to speculatively execute multiple future tokens in parallel with the current token using predicted values based on early-layer hidden states. FREE (Bae et al., 2023) proposes a shallow-deep module and a synchronized parallel decoding to improve the efficiency. EESD (Liu et al., 2024) proposes an early-exiting framework with a self-distillation method and leverages Thompson Sampling to regulate the generation processes. For retrieval-based SD, Token Recycling (Luo et al., 2024) proposes a method to utilize the dropped draft tokens and generate draft tokens via an adjacency matrix. Different from our method, Token Recycling is a query-dependent method that utilizes information from other queries, which may result in limitations when applied to real-world applications with complex and dynamic user inputs. SAM Decoding (Hu et al., 2024) utilizes a common text corpus and dynamic text sequence as retrieved sources and proposes a suffix automaton to efficiently obtain more accurate match positions.
Appendix B Details of Motivation Experiments
To investigate the prediction ability of the last logit, we conduct two motivation experiments to demonstrate its effectiveness and robustness. For the effectiveness of the last logit, as shown in Figure 2, we conduct autoregressive inference for the 6 models and record the logits for each decoded token. Then, for each decoded token $x_{i}$ which is sampled from the logit $p_{i-1}$, we investigate the rank of $x_{i}$ in $p_{i-2}$, i.e., the corresponding last logit, and visualize the statistics of the rank in Figure 2. For the robustness of the last logit, as shown in Figure 3, we conduct last logit decoding and investigate its mean accepted tokens per decoding step. Both experiments are conducted on a small subset of Spec-Bench, where we randomly sample 2 questions for each sub-category of MT, and 10 questions for other categories (Trans, Sum, QA, Math, RAG).
Why Not Use Next-Next-Next Tokens (NNNT)?
A natural question is whether we can extend the approach to predict even further tokens. We investigated this: the ground-truth NNNT appears in the top-60 last logits with $\geq$40% probability, showing the phenomenon does extend.
However, we did not use NNNT due to combinatorial explosion. If we speculate top-60 for both NNT and NNNT, the draft tree would contain at least $60^{2}=3{,}600$ tokens, far exceeding efficient verification capacity. As shown in Table 15 (Appendix E.8), tree capacity beyond 64-128 tokens yields diminishing returns as verification overhead dominates.
Our chosen approach (NNT + retrieval with K=64) represents the optimal efficiency-accuracy trade-off given current hardware constraints.
Intuition: Why Can Last Logits Predict Next-Next Tokens?
While our contribution is an empirical discovery, we offer intuition for why this phenomenon exists:
The last logit $p_{i-1}=P(x_{i}|x_{1},\ldots,x_{i-1})$ is computed from hidden states encoding the full context up to position i1i-1. While the model is trained to predict $x_{i}$, the rich contextual representation necessarily contains information about likely continuations beyond $x_{i}$—otherwise, the model could not generate coherent multi-token sequences.
Our findings suggest that top-ranked tokens in $p_{i-1}$ correlate with likely values of $x_{i+1}$, particularly in structured text where multi-token patterns (phrases, named entities, common collocations) are prevalent. This predictive signal exists without fine-tuning, making it immediately exploitable for inference acceleration.
We leave formal theoretical analysis of this phenomenon to future work.
Appendix C Pseudo Code to Prepare Attention Inputs
We provide pseudo code to organize the retrieved multiple draft token sequences into a draft tree and prepare its attention mask. More detailed implementations can be found in our attached code.
def prepare_attention_inputs(past_len, next_token, candidate_list, num_draft_tokens):
’LogitSpec organizes draft tokens in a tree manner. Each sub-sequence corresponds to a local causal mask.’
seq_len = num_draft_tokens + 1
# organize the candidate list into a sequence
draft_ids = [next_token] + [token for sub in candidate_list for token in sub]
# prepare original position ids and attention mask
position_ids = torch.zeros((1, seq_len), dtype=torch.long)
causal_mask = torch.full((seq_len, past_len + seq_len), fill_value=0)
causal_mask[:, :past_len+1] = 1
# prepare causal mask
idx = 1
for sub_sequence in candidate_list:
l = len(sub_sequence)
sub_mask = torch.tril(torch.ones((l, l)))
causal_mask[idx:idx+l, idx+past_len:idx+past_len+l] = sub_mask
position_ids[0, idx:idx+l] = torch.arange(l) + 1
idx += l
position_ids += past_len
return draft_ids, causal_mask, position_ids
Appendix D More Implementation Details
D.1 Baselines
We compare LogitSpec against strong speculative decoding baselines. For draft-model-free, plug-and-play methods, we include: (a) Lookahead Decoding (Fu et al., 2024), which utilizes dropped draft tokens as a retrieval corpus; (b) REST (He et al., 2024), which employs an external knowledge base as the retrieval model; and (c) PLD (Saxena, 2023), which extracts n-grams from the current user-input prompt and decoded tokens as the retrieval model. In addition, we report results for a draft-model-based baseline, vanilla speculative sampling (SpS) (Leviathan et al., 2023; Chen et al., 2023) with a small Vicuna-68M draft model, to illustrate the trade-off between extra model parameters / training and acceleration. We use Vicuna series models (7B, 13B, 33B), Llama-2 series models (7B, 13B, 70B), Llama-3.1-Instruct-8B and Qwen-3-8B as backbones, and report results with baseline methods on Vicuna and Llama-2 as the main results, where the performance of baseline methods can be reproduced with their official implementations.
D.2 Hardware and Software Configurations
We follow the setting of Spec-Bench with Pytorch (Paszke et al., 2019) 2.6.0 and CUDA (NVIDIA et al., 2020) 12.4. Our implementation is based on Hugging Face transformers. For main experiments with Vicuna and Llama-2 models (Section 5.1 and all baseline comparisons), we use transformers (Wolf et al., 2020) version 4.37.1. For additional experiments with Llama-3.1-8B-Instruct and Qwen3-8B (Section E.1), we use transformers version 4.57.1, which is necessary to support these newer model architectures. Experiments with models between 7B and 33B are conducted on a single NVIDIA SXM A100 80G GPU, while experiments with 70B models are conducted on 2 NVIDIA SXM A100 80G GPUs. The inference precision is float16. Following prior work, we conduct experiments with temperature 0 and batch size 1. The maximal generation lengths are 1024.
D.3 Retrieval Process
The retrieval process significantly affects overall acceleration. A naive implementation would be computationally prohibitive for LogitSpec.
Naive Approach.
For each of $k$ speculated next-next tokens, scan the entire prompt (length $L$) to find matching $m$-grams, yielding $O(k\cdot m\cdot L)$ complexity.
Our Hash-Based Optimization.
We reduce this to $O(n+k)$ per decoding step:
  1. 1.
    Preprocessing (amortized): Slide a window of size $m$ over the prompt once, building a hash table mapping each $m$-gram (key) to its continuation span (up to $n$ tokens as value). Cost: $O(L)$, done once per prompt.
  2. 2.
    Per-step retrieval: For each of $k$ next-next token candidates:
    • Construct the $m$-gram query: $O(m)\approx O(1)$ (since $m$ is small, typically 3)
    • Hash table lookup: $O(1)$
    • Copy continuation span: $O(n)$
    Total per step: $O(k\cdot(1+n))=O(k+kn)$.
Memory Cost.
The hash table requires $O(mL)$ memory, which is negligible compared to KV cache (typically several GB for long contexts).
Empirical Validation.
Figure 5 shows retrieval accounts for only 1.17% of total wall-clock time, confirming our optimization is effective in practice.
D.4 Evaluation Instructions
In our experiments, we employ different instructions for different evaluation tasks. Specifically, for Spec-Bench, we use its standard instructions:
Prompt Templates for Spec-Bench A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: Question
ASSISTANT:
For CNN/DM, we prepend “Summarize: ” to the question:
Prompt Templates for CNN/DM A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: Summarize: Question
ASSISTANT:
For GSM8K, we follow the setting with (Liu et al., 2025) and use an 8-shot CoT for inference:
Prompt Templates for GSM8K A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: 8-shot CoT Q: Question
ASSISTANT:
For HumanEval, we add a simple instruction “Please help me to complete this code, just output your codes directly.”:
Prompt Templates for HumanEval A chat between a curious user and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the user’s questions.
USER: Please help me to complete this code, just output your codes directly. Question
ASSISTANT:
D.5 Dataset Configurations
In our experiments, we evaluate LogitSpec on four categories of text generation tasks, including Spec-Bench, CNN/DM, GSM8K, and HumanEval. For Spec-Bench and HumanEval, we use the full data for evaluation. For CNN/DM and GSM8K, we randomly sample 1000 questions for evaluation. The maximal generation length is set as 1024 across all the experiments.
D.6 Model Configurations
In our experiments, all the models are deployed with precision float16. The <|eos token|> matches the tokenizer’s <|eos token|>. To effectively alleviate the overhead of rolling back the KV cache to the accepted draft tokens, we follow Medusa (Cai et al., 2024) to allocate continuous GPU memory for all the KV cache.
Ethics Statement
This work adheres to the ACL Code of Ethics. Our study does not involve human subjects or personally identifiable information, and we only use publicly available datasets under their respective licenses. We transparently report our methods and potential risks and do not recommend deployment in high-stakes settings without further safety assessments.
Appendix E More In-depth Analysis of LogitSpec
E.1 More Results on Different LLM Backbones
Comparison with State-of-the-Art Retrieval-Based Methods.
To position LogitSpec against recent advances in retrieval-based speculative decoding, we compare with SAM Decoding (Hu et al., 2024) and Token Recycling (Luo et al., 2024). Both methods are plug-and-play and do not require draft models, making them directly comparable to LogitSpec.
[…]
  • vs. Token Recycling: Token Recycling reuses previously rejected draft tokens via an adjacency matrix, which is query-dependent and requires cross-query information. In contrast, LogitSpec is query-independent and uses the target model’s own predictive capability to guide retrieval, making it more robust when query patterns are diverse.
  • vs. SAM Decoding: SAM uses suffix automata for efficient n-gram matching from a static corpus. While this improves matching speed, it still faces the fundamental challenge of low match rates when context repetition is limited. LogitSpec addresses this by speculating next-next tokens, effectively expanding the retrieval space beyond observed n-grams.
[…]
Here, we provide additional results for different LLM backbones to provide more insights into LogitSpec, including Llama 2 chat series models, LLaMA-3.1-8B-Instruct, and Qwen3-8B.
We present the results of Llama 2 chat series models on Spec-Bench, CNN/DM, GSM8K, and HumanEval in Table 7 and Table 8. LogitSpec consistently achieves the best acceleration among all baselines across Spec-Bench, CNN/DM, GSM8K, and HumanEval, further validating its effectiveness. For completeness, Table 5 reports the per-subtask Spec-Bench breakdown on Vicuna 7B/13B/33B, complementing the summary statistics in the main text.
E.2 Per-Task Spec-Bench Breakdown on Vicuna
Per-task Analysis.
The detailed breakdown reveals important insights into LogitSpec’s robustness across diverse tasks. In the Translation task, where almost no reference is available to serve as draft tokens, PLD and Lookahead achieve poor acceleration. Even equipped with a large external database ($\approx$ 12 GB), REST only achieves $1.17\times\sim 1.31\times$ speedup, while LogitSpec achieves $1.38\times\sim 1.43\times$ speedup without external databases, demonstrating the prediction ability of the last logit on the next next token. In contrast, on tasks with high context repetition such as Summarization and RAG, LogitSpec achieves even stronger performance, highlighting how next-next token speculation complements retrieval-based approaches across varying task characteristics.
Table 5: Experimental results of LogitSpec on Spec-Bench (Xia et al., 2024) with Vicuna. We report the speedup ratio on each sub task, mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
Models Method MT Trans Sum QA Math RAG MAT Speedup
Vicuna 7B Lookahead 1.40 1.14 1.19 1.24 1.55 1.09 1.66 1.27
REST 1.63 1.31 1.36 1.66 1.21 1.73 1.82 1.48
PLD 1.64 1.04 2.43 1.14 1.61 1.71 1.73 1.59
SpS 1.66 1.13 1.62 1.49 1.47 1.55 2.28 1.49
LogitSpec 1.92 1.39 2.68 1.70 2.22 1.87 2.44 2.01
Vicuna 13B Lookahead 1.30 1.06 1.20 1.12 1.48 1.12 1.63 1.22
REST 1.52 1.17 1.37 1.53 1.19 1.55 1.82 1.38
PLD 1.47 1.02 2.19 1.03 1.57 1.71 1.68 1.48
SpS 1.60 1.13 1.68 1.39 1.53 1.67 2.18 1.49
LogitSpec 1.89 1.43 2.33 1.43 2.23 1.93 2.32 1.93
Vicuna 33B Lookahead 1.32 1.08 1.20 1.06 1.54 1.15 1.61 1.24
REST 1.63 1.27 1.45 1.61 1.30 1.61 1.80 1.48
PLD 1.44 1.06 2.00 1.07 1.55 1.45 1.55 1.42
SpS 1.75 1.28 1.76 1.53 1.69 1.68 2.01 1.61
LogitSpec 1.77 1.38 2.15 1.37 2.00 1.69 2.13 1.76
Table 6 compares LogitSpec against SAM Decoding and Token Recycling on AIME, LongBench, and MATH with the Llama-3.1-8B-Instruct backbone under identical decoding settings ($T=0$, $K=64$, greedy sampling). LogitSpec consistently attains the best MAT and speedup across all datasets.
Table 6: Comparison with stronger retrieval-based baselines on Llama-3.1-8B-Instruct. We report MAT and end-to-end speedup on AIME, LongBench, and MATH, as well as the macro average. All methods share the same hardware, decoding parameters (temperature $T=0$), and tree capacity $K=64$.
Method AIME LongBench MATH Average
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
[…]
We also present more results applying LLaMA-3.1-8B-Instruct and Qwen3-8B on Spec-Bench, CNN/DM, GSM8K, and HumanEval in Tables 9 and 10. We observe that compared to standard autoregressive decoding, LogitSpec consistently achieves around $2\times$ inference acceleration on average across these datasets.
Table 7: Experimental results of LogitSpec on CNN/DM (Nallapati et al., 2016), GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) with Llama-2. We report the mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
Models Method CNN/DM GSM8K HumanEval Overall
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Llama 2 7B Lookahead 1.58 1.36 2.02 1.67 1.77 1.53 1.79 1.52
REST 1.71 1.33 1.51 1.18 1.97 1.52 1.73 1.34
PLD 1.89 1.73 3.32 2.98 1.57 1.41 2.26 2.04
SpS 1.99 1.46 2.83 1.87 2.13 1.51 2.32 1.61
LogitSpec 2.41 2.02 4.44 3.68 2.17 1.75 3.01 2.48
Llama 2 13B Lookahead 1.56 1.18 2.08 1.52 1.84 1.66 1.83 1.45
REST 1.71 1.34 1.53 1.26 1.96 1.63 1.73 1.41
PLD 1.89 1.52 3.24 2.54 1.73 1.63 2.29 1.90
SpS 1.95 1.34 2.87 1.85 2.33 1.76 2.38 1.65
LogitSpec 2.43 2.03 4.31 3.24 2.38 2.10 3.04 2.46
Llama 2 70B Lookahead 1.53 1.28 1.90 1.57 1.86 1.57 1.76 1.47
REST 1.67 1.35 1.63 1.32 1.96 1.66 1.75 1.44
PLD 1.98 1.74 1.63 1.46 1.62 1.49 1.74 1.56
SpS 2.01 1.71 1.98 1.69 2.21 1.70 2.07 1.70
LogitSpec 2.67 2.10 2.37 1.87 2.33 1.93 2.46 1.97
Table 8: Experimental results of LogitSpec on Spec-Bench (Xia et al., 2024) with Llama-2. We report the speedup ratio on each sub task, mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results and underline the suboptimal results for each backbone model.
Models Method MT Trans Sum QA Math RAG MAT Speedup
Llama 2 7B Lookahead 1.55 1.44 1.42 1.48 1.69 1.46 1.69 1.51
REST 1.58 1.20 1.38 1.61 1.31 1.55 1.83 1.48
PLD 1.46 1.34 1.89 1.23 1.65 1.58 1.59 1.49
SpS 1.57 1.38 1.55 1.46 1.55 1.60 2.07 1.53
LogitSpec 1.83 1.72 2.19 1.63 2.15 1.94 2.15 1.87
Llama 2 13B Lookahead 1.39 1.36 1.21 1.39 1.69 1.25 1.68 1.37
REST 1.53 1.14 1.30 1.51 1.23 1.45 1.85 1.42
PLD 1.36 1.19 1.58 1.16 1.62 1.37 1.56 1.37
SpS 1.52 1.26 1.43 1.42 1.54 1.50 2.03 1.48
LogitSpec 1.73 1.57 1.86 1.53 2.14 1.73 2.12 1.74
Llama 2 70B Lookahead 1.45 1.35 1.28 1.38 1.71 1.31 1.66 1.41
REST 1.63 1.33 1.38 1.67 1.35 1.55 1.83 1.53
PLD 1.34 1.32 1.76 1.18 1.63 1.47 1.51 1.39
SpS 1.65 1.50 1.62 1.57 1.70 1.68 1.88 1.63
LogitSpec 1.66 1.58 1.95 1.58 2.03 1.78 2.12 1.72
Table 9: Experimental results of LogitSpec on Spec-Bench (Xia et al., 2024) with LLaMA-3.1-8B-Instruct and Qwen3-8B. We report the speedup ratio on each sub task, mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results for each backbone model.
Model Method MT Trans Sum QA MATH RAG MAT Speedup
Llama-3.1-8B-Instruct Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Llama-3.1-8B-Instruct LogitSpec 1.89 1.67 1.94 1.68 2.01 1.77 2.11 1.88
Qwen-3-8B Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Qwen-3-8B LogitSpec 1.71 1.74 1.64 1.65 1.89 1.68 1.95 1.75
Table 10: Experimental results of LogitSpec on CNN/DM (Nallapati et al., 2016), GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021) with LLaMA-3.1-8B-Instruct and Qwen3-8B. We report the mean accepted tokens per decoding step (MAT) and overall speedup ratio. We bold the best results for each backbone model.
Model Method CNN/DM GSM8K Humaneval Overall
MAT Speedup MAT Speedup MAT Speedup MAT Speedup
Llama-3.1-8B-Instruct Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Llama-3.1-8B-Instruct LogitSpec 2.04 1.85 2.18 1.95 2.63 2.31 2.28 2.04
Qwen-3-8B Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Qwen-3-8B LogitSpec 1.77 1.59 2.18 1.94 2.18 1.92 2.04 1.82
E.3 More Results on Different Benchmarks
Table 11: Results of LogitSpec on AIME datasets.
Method AIME24 AIME25
MAT Speedup MAT Speedup
Vanilla 1.00 1.00 1.00 1.00
LogitSpec 3.41 3.25 3.76 3.33
We provide more results for LogitSpec on complex math reasoning tasks including MATH (Hendrycks et al., 2021) and AIME 24 & 25 (Mathematical Association of America, 2025) datasets in Figure 6 and Table 11. For each of these benchmarks, we report the real-world speedup for each subset, the overall mean accepted tokens per decoding step (MAT), and the overall speedup. These expanded experimental results further corroborate LogitSpec’s effectiveness and applicability, yielding an overall MAT of 3.32 and an overall speedup of $2.78\times$ on MATH, and up to 3.76 MAT and $3.33\times$ speedup on AIME25. This demonstrates LogitSpec’s robust performance even on challenging reasoning tasks.
Figure 6: LogitSpec performance on MATH benchmark subsets.
E.4 Sampling with Non-Zero Temperatures
Why LogitSpec Remains Effective Under Sampling.
A key question is whether LogitSpec’s reliance on top-$k$ logits makes it sensitive to temperature scaling. We clarify the mechanism and provide empirical validation.
Candidate Selection is Temperature-Invariant. In our implementation, speculative next-next token candidates are selected from the raw logits before applying temperature. Since temperature scaling is a monotonic transformation, it preserves ranking:
$z_{i}>z_{j}\quad$ $\Longleftrightarrow\quad\exp(z_{i}/T)>\exp(z_{j}/T)$
$\Longleftrightarrow\quad p_{i}(T)>p_{j}(T),\quad\forall T>0$ (9)
where $p(T)=\text{softmax}(z/T)$. Therefore, the top-$k$ candidate set used for speculation is identical across all temperatures and does not “fill with noise tokens” as $T$ increases.
What Changes: Acceptance Probability. Higher temperatures make the actual sampled trajectory deviate more from the most-likely path, reducing acceptance rates. This is an inherent property of all speculative decoding methods under sampling, not specific to LogitSpec.
To study the robustness of LogitSpec under stochastic decoding, we evaluate performance across temperatures $T\in\{0.2,0.4,0.6,0.8\}$ on AIME, LongBench, and MATH with Llama-3.1-8B-Instruct. As shown in Table 12, LogitSpec maintains strong performance even at higher temperatures, still achieving speedups above $1.8\times$ even at $T=0.8$. This demonstrates that LogitSpec remains effective across diverse generation scenarios.
Table 12: LogitSpec under different sampling temperatures on Llama-3.1-8B-Instruct. Draft candidates are drawn from raw logits before temperature scaling, so the speculative candidate set is unchanged across $T$; higher $T$ only changes the sampled path.
Temperature AIME MAT AIME Speedup LongBench MAT LongBench Speedup MATH MAT MATH Speedup
[…]
E.5 Cross-Lingual Next-Next Token Coverage
To validate the universality of the last logit’s predictive capability across different linguistic structures, we conduct experiments on the MMMLU benchmark covering five languages: English, Chinese, Japanese, German, and French. Table 13 reports the fraction of ground-truth next-next tokens appearing in the top-$k$ logits for each language. The results show consistent coverage across all languages: at $k=64$, over $50\%$ of next-next tokens fall within the top-$k$ predictions for every language. This demonstrates that our observation about the last logit’s predictive capability is not restricted to English or specific linguistic structures, but rather represents a fundamental property of modern LLMs that generalizes across diverse languages with different word orders and morphological characteristics.
Table 13: Fraction of ground-truth next-next tokens that appear within the top-$k$ logits of the last token on MMMLU subsets (Llama-3.1-8B-Instruct). Each language averages 20 questions.
Language $k=4$ $k=8$ $k=16$ $k=32$ $k=64$
English 15.25% 22.02% 33.60% 42.06% 51.83%
Chinese 17.14% 24.21% 36.29% 45.68% 55.07%
Japanese 15.16% 24.30% 34.72% 46.17% 55.87%
German 14.64% 21.44% 32.87% 42.42% 50.64%
French 14.86% 21.77% 32.21% 42.78% 52.22%
These results validate a key claim from our rebuttal: the last logit’s predictive capability is not restricted to English or specific linguistic structures. The consistency across languages with markedly different characteristics—Chinese (logographic, subject-verb-object), Japanese (agglutinative, subject-object-verb), German (inflectional, flexible word order), and French (Romance, subject-verb-object)—suggests this is a fundamental property of modern LLMs’ internal representations rather than a language-specific artifact. Combined with our diverse task coverage (reasoning, QA, summarization, translation, code generation), this cross-lingual validation provides strong evidence for the generality of our core observation.
E.6 More Results on Long-Context Scenarios
[…]
Figure 7: LogitSpec performance on LongBench subsets.
E.7 Ablation Study on Different Pruning Strategies
In Section 4.2, we apply a heuristic pruning strategy to control the number of draft tokens. The role of the pruning algorithm is to effectively control the size of the resulting draft tree, keeping the decoding overhead low without sacrificing too many possibilities, which is a strategy to balance breadth and cost. To provide more insight into the pruning strategy, we conduct ablation studies in Table 14. Specifically, we consider two different pruning strategies:
  • (a)
    Strategy $1$: A rank-based heuristic: if the rank is <<44, we preserve $5$ tokens; <<88, 44 tokens; <<$16$, 33 tokens; <<32, 22 tokens; else $1$ token.
  • (b)
    Strategy $2$: A simple heuristic: preserving $4$ tokens for all speculated next next tokens.
The results in Table 14 demonstrate that while different pruning strategies have some effect on the final performance, the overall performance of LogitSpec is quite robust. The speedup remains stable at approximately $1.9\times$ across these different approaches.
We also would like to clarify that the core contribution of our work is next next token speculation using the last logit to guide and enhance the efficiency and accuracy of retrieval-based speculative decoding. The primary purpose of the pruning algorithm is to serve as an auxiliary module for our core mechanism.
Table 14: Ablation study on pruning strategies. LogitSpec $_{with\hskip 2.84526pts1}$ and LogitSpec $_{with\hskip 2.84526pts2}$ denote LogitSpec with Strategy 1 and Strategy 2, respectively.
Method MT Trans Sum QA MATH RAG MAT Speedup
LogitSpec $_{with\hskip 2.84526pts1}$ 1.90 1.68 1.94 1.66 2.04 1.78 2.12 1.88
LogitSpec $_{with\hskip 2.84526pts2}$ 1.86 1.67 1.87 1.68 2.03 1.74 2.02 1.85
LogitSpec (ori) 1.89 1.67 1.94 1.68 2.01 1.77 2.11 1.88
E.8 Ablation Study on Draft Tree Capacity
As mentioned in Section 4.2, we set the capacity of the draft tree to K=64. To provide further insight, we evaluate a range of $K$ values from $32$ to $128$ in Table 15. These results reveal a clear trade-off: as K increases, the MAT per step improves (from $2.03$ to $2.30$), since a larger tree offers more opportunities for token acceptance. However, the overall speedup ratio peaks at K=64 ($1.92\times$) and subsequently declines. This is because verifying an overly large draft tree introduces significant computational overhead that ultimately negates the advantages of a higher acceptance rate.
Table 15: Ablation study on pruning hyperparameters $K$ using LLaMA-3.1-8B-Instruct as the backbone model.
Method MT Trans Sum QA MATH RAG MAT Speedup
Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
LogitSpec 32 1.79 1.62 1.82 1.56 1.93 1.67 1.95 1.79
LogitSpec 48 1.86 1.66 1.90 1.67 2.00 1.73 2.04 1.85
LogitSpec 64 1.89 1.67 1.94 1.68 2.01 1.77 2.11 1.88
LogitSpec 80 1.85 1.64 1.86 1.67 1.84 1.74 2.14 1.83
LogitSpec 96 1.81 1.65 1.83 1.66 1.84 1.73 2.17 1.80
LogitSpec 112 1.79 1.61 1.77 1.64 1.83 1.69 2.21 1.77
LogitSpec 128 1.74 1.62 1.73 1.58 1.81 1.62 2.22 1.73
E.9 Real-World Example
To provide more insights into LogitSpec, we present a real-world example to illustrate how “next next token speculation” allows LogitSpec to succeed where standard retrieval methods fail.
Taking the prefix
Q: A pen costs as much as a pencil and eraser combined. A pencil costs \$1.20 and an eraser costs \$0.30. How much will 8 pens cost?
A: To find the cost of 8 pens, we first need to find the cost of one pen. A pencil costs \$1.20 and an eraser costs \$0.30. The
At the first step, LogitSpec generates the next token “combined” and then speculates the next next token with the following high-probability candidates:
  • total
  • cost
  • combination
  • pen
  • eraser
Then, LogitSpec extends each candidate with retrieved n-grams and verifies these draft sequences:
  • $\times$] . A pencil costs
  • $\times$] total they had
  • $\checkmark$] cost of one pen
  • $\times$] combination (no matched n-grams retrieved)
  • $\times$] pen.
  • $\times$] eraser costs \$
Finally, with the guidance of the last logit, we successfully accept $3$ draft tokens and generate “the combined cost of one pen”. In the next decoding step, LogitSpec again accepts $3$ draft tokens and generates “the combined cost of one pencil and one eraser”. However, without last-logit speculation, standard retrieval can only retrieve the first n-gram “A pencil costs”.
Appendix F Discussion: Integration with Production Frameworks
vLLM Integration Feasibility.
While our current implementation uses Hugging Face transformers, several reviewers asked about integration with production inference frameworks like vLLM. We discuss technical feasibility:
Tree Attention Support. vLLM has recently added native tree attention support:
Compatibility with vLLM’s Architecture. LogitSpec’s draft tree can be naturally represented as multiple sequences with shared prefixes, aligning with vLLM’s batched inference paradigm. Each branch corresponds to a different draft sequence, similar to the prefill process for multiple sequences. The nano-PEARL repository demonstrates a similar integration approach for tree-based verification in vLLM.
Minimal Overhead. Beyond tree attention, LogitSpec only requires:
  • The last logit at each step (already computed by vLLM)
  • CPU-based retrieval and tree construction (can run in parallel with GPU inference)
We conclude that there are no fundamental algorithmic obstacles to deploying LogitSpec in vLLM; the remaining work is engineering integration, which we leave to future work.
Appendix G LLM Usage
We used a large language model (LLM)–based writing assistant solely for grammar and wording improvements on draft text. The LLM did not generate research ideas, claims, proofs, algorithms, code, figures, or analyses, and it did not have access to any non-public data. All edits suggested by the LLM were manually reviewed and either accepted or rewritten by the authors, who take full responsibility for the final content. The LLM is not an author of this paper.